14. Consider a datagram network using 8-bit host addresses. Suppose a router use longest prefix matching and has the following forwarding table Interface Prefix Match 01 0 10 001 2 For each of the four interfaces, give the associated range of destination host addresses and the number of addresses in the range.
Answer #1

Destination Address Range Link Interface Range

01000000 through 01111111 0 2^6 = 64

00000000 through 00011111 1 2^5 = 32

10000000 through 10111111 2 2^6 = 64

00100000 through 00111111 2 2^5 = 32

11000000 through 11111111 3 2^6 = 64

number of addresses in each range = 2^N , where N = (Number of maxmimum bit in the host address - number of bits in prefix match).

Here for interface 0, 8-2 =6. Hence 2^6 = 2*2*2*2*2*2 = 64.

For interface 1, 8-3 = 5. Hence 2^5 = 2*2*2*2*2 = 32.